VELOCITY-MAPPED ION IMAGING OF METHYL NITRITE PHOTODISSOCIATION

\underline{NICLAS A. WEST}, AIDAN J. KLOBUCHAR, JEFFREY A. BARTZ, Department of Chemistry, Kalamazoo College, 1200 Academy Street, Kalamazoo, MI 49006.

Methyl nitrite, CH_{3}ONO, was photodissociated at 355 nm in a velocity-map ion imaging apparatus. The NO photoproduct was detected through state-resolved 1+1’ resonance-enhanced multiphoton ionization. The NO products are highly oriented in the laboratory frame, indicating perpendicular transition and a prompt dissociation. The velocities of the NO products reveal that methyl nitrite undergoes rapid predissociation after 355 nm excitation.